# Capacity Gains in MIMO Systems with Few-Bit ADCs Using Nonlinear Analog Circuits

### v2

@@ IV-B2 Low-resolution ADCs and Concatenated Sequences of Envelope Detectors
Proposition 4 is an extension of the properties shown in the proof of Theorem 1. We provide a brief justification of each property in the following. Property 1 follows by the fact that the number of codewords in $\mathcal{C}$ is equal to the number of roots of $f_{j,k},j\in[n_{q}],k\in[\ell-1]$ plus one (e.g., see Figure 6). Property 2 and 6 follow by a similar argument as Equation (20) in proof of Proposition 1. Properties 3 and 5 follow by the fact that each root of $f_{j,k},j\in[n_{q}],k\in[\ell-1]$ corresponds to a value transition in the output of exactly one of the ADCs (since the roots are not repeated) and at each transition the value changes either one unit up or down since in the input crosses one threshold at a time at its value is changed continuously. Property 4 follows by the fact that the transition count at each position is equal to the number of roots of $f_{j,k},k\in[\ell-1]$ for a fixed $j\in[n_{q}]$.
As a step towards characterizing capacity when $\ell>2,\delta>1$, we first study the capacity region for systems with one-bit ADCs, i.e., $\ell=2,\delta>1$. To this end, we prove two useful propositions. The first one shows that given an ordered set $\mathcal{C}$ satisfying the properties in Proposition 4 and a sequence of real numbers $(r_{1},r_{2},\cdots,r_{\gamma})$ satisfying the properties in Proposition 3, one can always construct a quantizer whose associated code is equal to $\mathcal{C}$ and whose roots sequence is $(r_{1},r_{2},\cdots,r_{\gamma})$. The second proposition provides conditions under which there exists a code satisfying the properties in Proposition 4. The proof ideas follow techniques used in study of balanced and locally balanced gray codes [50, 51]. Combining the two results allows us to characterize the necessary and sufficient conditions for existence of quantizers with desirable properties.
In the statement of the following proposition, for a given code $\mathcal{C}$, we have used the notation $\xi_{1},\xi_{2},\cdots,\xi_{\gamma-1}$ for the transition sequence of $\mathcal{C}$. That is, $\xi_{k},k\in\{1,\dots,\gamma-1\}$ is the bit position which is different between $\mathbf{c}_{k-1}$ and $\mathbf{c}_{k}$. We have defined the transition sets $\mathcal{I}_{j}\triangleq\{s|\xi_{s}=j\},j\in[n_{q}]$. Note from Property 5) in Proposition 4, we have $|\mathcal{I}_{j}|=\kappa_{j}=2^{\delta}$.
@@
Let $\ell=2,n_{q}\in\mathbb{N}$ and $\delta>1$ and consider an ordered set $\mathcal{C}\subset\{0,1\}^{n_{q}}$ satisfying properties 1)-5) in Proposition 4, and a sequence of increasing real numbers $r_{1},r_{2},\cdots,r_{\gamma}$, where $\gamma=n_{q}2^{\delta}$, such that $(r_{i_{s}},s\in\mathcal{I}_{j})$ is fully-symmetric for all $j\in[n_{q}]$, where $\mathcal{I}_{j}$ are the transition sets of $\mathcal{C}$. Then, there exists a quantizer $Q(\cdot)$ with associated analog functions $f_{j}(\cdot),j\in[n_{q}]$ such that its associated code is $\mathcal{C}$, and $r_{1},r_{2},\cdots,r_{\gamma}$ is the non-decreasing sequence of roots of its associated analog functions $f_{j,k}(\cdot),j\in[n_{q}],k\in[\ell-1]$.
Proof.
Please see Appendix C. ∎
@@
(Code Construction) Let $\ell=2$, $n_{q}\in\mathbb{N}$, and $\kappa_{1}$, $\kappa_{2},\cdots,\kappa_{n_{q}}$ be even numbers such that $|\kappa_{j}-\kappa_{j^{\prime}}|\leq 2,j,j^{\prime}\in[n_{q}]$. Then, there exists a code $\mathcal{C}$ with transition count at position j equal to $\kappa_{j},j\in[n_{q}]$ satisfying properties 1), 2), 3), and 5) in Proposition 4 such that $|\mathcal{C}|=\min\{2^{n_{q}},\sum_{j=1}^{n_{q}}\kappa_{j}\}$. Particularly, if $\kappa_{j}=2^{\delta},j\in[n_{q}]$, then there exists $\mathcal{C}$ with $|\mathcal{C}|=\min\{2^{n_{q}},n_{q}2^{\delta}\}$ satisfying properties 1)-5) in Proposition 4.
Proof.
Please see Appendix D. ∎
Using Propositions 5 and 6, we characterize the channel capacity for $\ell=2$ and $\delta>1$. Let us define $\Gamma_{2}(n_{q},\ell)\triangleq\min(2^{n_{q}},n_{q}2^{\delta}+1)$ and the set $\mathcal{T}_{\Gamma_{2}}\subseteq\mathbb{R}^{\Gamma_{2}-1}$ as the set of sequences of increasing real numbers $r_{1},r_{2},\cdots,r_{\gamma}$, where $\gamma=n_{q}2^{\delta}$ for which there exists a partition $\{\mathcal{I}_{j},j\in[n_{q}]\}$ of $[\delta]$ such that $(r_{i_{s}},s\in\mathcal{I}_{j})$ is fully-symmetric for all $j\in[n_{q}]$, and there exists a code satisfying Properties 1)-5) in Proposition 4 whose transition sets are equal to $\mathcal{I}_{j},j\in[n_{q}]$ and which has exactly one repeated codeword, i.e., only the first and last codewords are repeated. The following theorem characterizes the channel capacity.
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\in\mathbb{R}$, $\delta>1$, and assume that the ADC resolution is restricted to $\ell=2$. Then, the capacity is given by:
$C_{env}(P_{T},P_{ADC},h,\delta)=\max_{\begin{subarray}{c}\gamma=\Gamma_{2}(n_{q},\ell):\\ (n_{q},\ell)\in\mathcal{N}(P_{ADC})\end{subarray}}\sup_{\mathbf{x}\in\mathbb{R}^{\gamma}}\sup_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\sup_{\mathbf{t}\in\mathcal{T}_{\gamma}}I(X;\widehat{Y}),$ (6)
where $\widehat{Y}=Q(hX+N)$, $N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of distributions on $\{x_{1},x_{2},\cdots,x_{\gamma}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $Q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,\gamma-1\}$ and $Q(y)=0$ if $y>t_{\gamma-1}$ or $y<t_{1}$.
The proof follows by similar arguments as in the proof of Theorem 1. The converse follows from Proposition 4 Item 4). Achievability follows from Proposition 6.
The region given in Theorem 2 is difficult to analyze since finding the set $\mathcal{T}_{\gamma}$ may be computationally complex. Inner bounds to the achievable region may be numerically derived by assuming additional symmetry restriction such as uniform quantization restrictions. This is studied in more detail in the numerical evaluations provided in Section VII.
For scenarios with $\ell>2$ and $\delta>1$, let us define $\Gamma_{\ell}\triangleq\min(\ell^{n_{q}},(\ell-1)n_{q}\ell^{\delta}+1)$ and the set $\mathcal{T}_{\Gamma_{\ell}}\subseteq\mathbb{R}^{\Gamma_{\ell}-1}$ as the set of sequences of increasing real numbers $r_{1},r_{2},\cdots,r_{\gamma}$ satisfying the properties in Proposition 3, for which there exists a code $\mathcal{C}$ satisfying Properties 1)-5) in Proposition 4 such that i) the sets $\mathcal{P}_{j,k},j\in[n_{q}],k\in[\ell-1]$ in 3 are the indices of the codewords of $\mathcal{C}$ which have transition to or from value $k$ in their $j$th element, and ii) $\mathcal{C}$ has exactly one repeated codeword, i.e., only the first and last codewords are repeated. The following theorem characterizes the channel capacity. The proof follows from Propositions 4 and 6 similar to the arguments given in the proof of Theorem 1.
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\in\mathbb{R}$, and $\delta\in\mathbb{N}$. Let Γ,subscriptnormal-$\Gamma_{\ell},\ell\in\mathbb{N}$ be the maximum size of codes satisfying condition 1)-5) in Proposition 4. Then,
$C_{env}(P_{T},P_{ADC},h,\delta)=\max_{\begin{subarray}{c}\gamma=\Gamma_{\ell}(n_{q},\ell):\\ (n_{q},\ell)\in\mathcal{N}(P_{ADC})\end{subarray}}\sup_{\mathbf{x}\in\mathbb{R}^{\gamma}}\sup_{P_{X}\in\mathcal{P}_{\mathbf{x}}}\sup_{\mathbf{t}\in\mathcal{T}_{\gamma}}I(X;\widehat{Y}),$ (9)
where $\widehat{Y}=Q(hX+N),N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ consists of distributions on $\{x_{1},x_{2},\cdots,x_{\gamma}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $Q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in[\gamma-1]$ and $Q(y)=0$ if $y>t_{\gamma-1}$ or $y<t_{1}$.
Optimizing (9) requires calculating $\Gamma_{\ell}(n_{q},\ell)$. The total number of codes satisfying conditions 1)-5) in Proposition 4 is bounded from above by (${(\ell-1)2^{\delta}n_{q}\choose(\ell-1)2^{\delta},(\ell-1)2^{\delta},\cdots,(\ell-1)2^{\delta}}$. Hence, for systems with a few low-resolution ADCs and small $\delta$ ($\delta=1,2$), one can find $\Gamma_{\ell}(n_{q},\ell)$ by searching over all such codes.
@@ V SISO Systems with Polynomial Operators
In the previous section, we investigated the channel capacity for SISO systems (analog beamforming) when sequences of concatenated envelope detectors are used for analog signal processing. In this section, we evaluate the resulting channel capacity when analog polynomial operators are used instead of envelope detectors.
@@
Let $n_{q}=\delta=2$ and $\ell=3$ and consider a quantizer characterized by polynomials $f_{a,1}(y)=y^{2}+2y$ and $f_{a,2}(y)=y^{2}+3y,y\in\mathbb{R}$, and thresholds
$t(1,1)=3,\quad t(1,2)=0,\quad t(2,1)=10,\quad t(2,2)=18,$
@@
As a first step, we show the following proposition about properties of codes and their associated polynomial functions which is analogous to Proposition 4 which addressed codes and their associated envelope-detector-based analog processing functions. The proof follows by similar arguments as that of Proposition 4 and is omitted for brevity.
Proposition 7 (Properties of Code and Associated Polynomial Functions).
Consider a quantizer $Q(\cdot)$ with threshold matrix $\mathbf{t}\in\mathbb{R}^{n_{q}\times(\ell-1)}$ and associated polynomials $f_{j}(\cdot)\in\mathcal{F}^{\delta}_{poly},j\in n_{q}$, such that $f_{j,k}(\cdot)\triangleq f_{j}(\cdot)-t(j,k),j\in[n_{q}],k\in[\ell-1]$ do not have repeated roots. The associated code $\mathcal{C}$ satisfies the following:
  1. 1.
    The number of codewords in $\mathcal{C}$ is equal to $\gamma\triangleq(\ell-1)\delta n_{q}+1$, i.e. $\mathcal{C}=(\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{\gamma-1})$.
  2. 2.
    All elements of the first codeword $\mathbf{c}_{0}$ are either equal to $\ell-1$ or equal to 00, i.e. $c_{i,0}=0,i\in\{0,1,\cdots,\gamma-1\}$ or $c_{i,0}=\ell-1,i\in\{0,1,\cdots,\gamma-1\}$.
  3. 3.
  4. 4.
    The transition count at every position is $\kappa_{j}=\frac{\gamma}{n_{q}}=(\ell-1)\delta,j\in[n_{q}]$.
  5. 5.
    Let $i_{1},i_{2},\cdots,i_{\kappa}$ be the non-decreasingly ordered indices of codewords where the $j$th element has value-transitions. Then, the sequence $(c_{i_{1},j},c_{i_{2},j},\cdots,c_{i_{\kappa},j})$ is periodic, in each period it takes all values between 00 and $\ell-1$, and $|c_{i_{k},j}-c_{i_{k+1},j}|=1,k\in[\kappa-1]$ holds. Furthermore, $c_{i_{1},j}\in\{0,\ell-1\}$.
  6. 6.
    If $\delta$ is even, then $|\mathcal{C}|\leq min(\ell^{n_{q}},(\ell-1)\delta n_{q})$ and if $\delta$ is odd, then $|\mathcal{C}|\leq min(\ell^{n_{q}},(\ell-1)\delta n_{q}+1)$
Using Proposition 6 and Proposition 7, and following the arguments in the proof of Theorem 2, one can prove the following theorem.
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\in\mathbb{R}$, then for even-valued $\delta$, we have:
$C_{poly}(P_{T},P_{ADC},h,\delta)=\max_{\begin{subarray}{c}\gamma=\Gamma(n_{q},\ell):\\ (n_{q},\ell)\in\mathcal{N}(P_{ADC})\end{subarray}}\sup_{\mathbf{x}\in\mathbb{R}^{\gamma}}\sup_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\sup_{\mathbf{t}\in\mathbb{R}^{\gamma-1}}I(X;\widehat{Y}),$ (12)
@@
$\max_{\begin{subarray}{c}\gamma=\Gamma(n_{q},\ell):\\ (n_{q},\ell)\in\mathcal{N}(P_{ADC})\end{subarray}}\sup_{\mathbf{x}\in\mathbb{R}^{\gamma}}\sup_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\sup_{\mathbf{t}\in\mathbb{R}^{\gamma-1}}I(X;\widehat{Y})\leq C_{Q}(P_{T},P_{ADC},h,\delta)$ (15)
$\qquad\qquad\leq\max_{\begin{subarray}{c}\gamma^{\prime}=\Gamma^{\prime}(n_{q},\ell):\\ (n_{q},\ell)\in\mathcal{N}(P_{ADC})\end{subarray}}\sup_{\mathbf{x}\in\mathbb{R}^{\gamma^{\prime}}}\sup_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\sup_{\mathbf{t}\in\mathbb{R}^{\gamma^{\prime}-1}}I(X;\widehat{Y}),$ (18)
where $n_{q}=\frac{P_{ADC}}{2\alpha}$, $\Gamma(n_{q},\ell)\triangleq\min(\ell^{n_{q}},(\ell-1)\delta n_{q})$, $\Gamma^{\prime}(n_{q},\ell)\triangleq\min(\ell^{n_{q}},(\ell-1)\delta n_{q}+1)$ $\widehat{Y}=Q(hX+N)$, $N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of distributions on $\{x_{1},x_{2},\cdots,x_{\gamma}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $Q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,\gamma-1]$ and $Q(y)=0$ if $y>t_{\gamma-1}$ or $y<t_{1}$.
@@
1) It can be noted from Equations (6) and (15) that the capacity expression for odd and even values of $\delta$ are not the same. This is due to Property 6) in Proposition 7 which gives different number of unique codewords for odd and even $\delta$. The reason is that while even degree polynomials yield the same output sign as their input converges to $\infty$- ∞ and $\infty$, for odd degree polynomials the output signs are different as their input converges to $\infty$- ∞ and $\infty$. This can potentially yield a larger number of unique codewords in the associated code of the quantizer since the first and last codeword are not equal to each other. This is in contrast with Theorem 3, where the capacity expression is the same for even and odd values of $\delta$. The reason is that the for absolute values the output sign is positive as their input converges to $\infty$- ∞ and $\infty$.
2) The region given in Theorem 4 strictly contains that of Theorem 3 for the same value of $\gamma$. The reason is that envelope detectors generate absolute value functions which force a symmetric structure on the Voronoi regions of $Q(\cdot)$. This manifests in the fully-symmetric condition $\mathbf{t}\in\mathcal{T}_{\gamma}$ in Theorem 3 and the properties given in Proposition 3; whereas for polynomial functions, no such symmetry is required and hence the optimization in Theorem 4 is over all $\mathbf{t}\in\mathbb{R}^{\gamma}$. However, as shown in Section VIII generating polynomial operators of degree up to $\delta$ requires a larger power budget compared to concatenating $\delta$ envelope detectors. This points to a rate-power tradeoff in using envelope detectors and polynomials operators.
3) One potential approach to improve upon the capacity of the system in Theorem 3 is to augment the envelope detectors by linearly combining their output with the original signal. That is, to generate operators of the form $f(y)=|y-a|+by,a,b\in\mathbb{R}$ instead of $f(y)=|y-a|,a\in\mathbb{R}$. This removes the fully-symmetric condition $\mathbf{t}\in\mathcal{T}_{\gamma}$ in Theorem 3 and yields a larger channel capacity. However, such linear combinations are challenging to implement using analog circuits due to timing issues in synchronizing the output of the envelope detector with the original signal. We hope to address these implementation challenges in future works.
4) For envelope detectors, the dimension $\gamma$ is equal to $n_{q}(\ell-1)2^{\delta}$, whereas for polynomial operators, it is equal to $n_{q}(\ell-1)\delta$. So, the dimension of the optimization space increases faster when concatenating envelope detectors compared to when increasing the polynomial degree. That is, at high SNRs, the capacity in Theorem 3 is larger than that of Theorem 4 for the same value of $\delta>1$. This is also observed in the numerical evaluations in Section VII.

### v3

@@ IV-D Concatenated Envelope Detectors and Few-bit ADCs
Proposition 4 is an extension of the properties shown in the proof of Theorem 1. We provide a brief justification of each property in the following. Property 1 follows by the fact that the number of codewords in $\mathcal{C}$ is one more than the number of roots of $f_{j,k},j\in[n_{q}],k\in[\ell-1]$. Property 2 follows from the fact that as $y\to\pm\infty$, its absolute value becomes asymptotically large. Property 6 follows from the facts that i) $n_{q}$ ADCs with $\ell$ output levels cannot output more than $\ell^{n_{q}}$ distinct outputs, ii) the number of roots cannot exceed $(\ell-1)n_{q}2^{\delta}$, and iii) the first and last decision regions have the same ADC outputs. Properties 3 and 5 follow by the fact that each root of $f_{j,k},j\in[n_{q}],k\in[\ell-1]$ corresponds to a value transition in the output of exactly one of the ADCs (since the roots are not repeated) and at each transition the value changes either one unit up or down since in the input crosses one threshold at a time and its value is changed continuously. Property 4 follows by the fact that the transition count at each position is equal to the number of roots of $f_{j,k},k\in[\ell-1]$ for a fixed $j\in[n_{q}]$.
As a step towards characterizing capacity when $\ell>2,\delta>1$, we first study the capacity region for systems with one-bit ADCs, i.e., $\ell=2,\delta>1$. To this end, we prove two useful propositions. The first one shows that given an ordered set $\mathcal{C}$ satisfying the properties in Proposition 4 and a sequence of real numbers $(r_{1},r_{2},\cdots,r_{\gamma})$ satisfying the properties in Proposition 3, one can always construct a quantizer whose associated code is equal to $\mathcal{C}$ and whose roots sequence is $(r_{1},r_{2},\cdots,r_{\gamma})$. The second proposition provides conditions under which there exists a code satisfying the properties in Proposition 4. The proof follows by construction, using techniques in balanced and locally balanced Gray codes [38]. The constructive proof is detailed in [37]. Combining the two results allows us to characterize the necessary and sufficient conditions for existence of quantizers with desirable properties.
For a given code $\mathcal{C}$, we denote $\xi_{k},k\in[\gamma]$ as the bit position which is different between $\mathbf{c}_{k-1}$ and $\mathbf{c}_{k}$. We define the transition sets $\mathcal{I}_{j}\triangleq\{k|\xi_{k}=j\},j\in[n_{q}]$. Note from Property 5 in Proposition 4, we have $|\mathcal{I}_{j}|=\kappa_{j}=2^{\delta}$.
@@
Let $\ell=2,n_{q}\in\mathbb{N}$, and $\delta>1$ such that $n_{q}2^{\delta}\leq 2^{n_{q}}$, and consider an ordered set $\mathcal{C}\subset\{0,1\}^{n_{q}}$ satisfying properties 1-5 in Proposition 4, and a sequence of increasing real numbers $r_{1},r_{2},\cdots,r_{\gamma}$, where $\gamma=n_{q}2^{\delta}$, such that $(r_{i_{k}},k\in\mathcal{I}_{j})$ is fully-symmetric for all $j\in[n_{q}]$, where $\mathcal{I}_{j}$ are the transition sets of $\mathcal{C}$. Then, there exists a quantizer $\mathbf{q}(\cdot)$ with associated functions $f_{j}(\cdot)\in\mathcal{F}^{\delta}_{env},j\in[n_{q}]$ such that its code is $\mathcal{C}$, and $r_{1},r_{2},\cdots,r_{\gamma}$ is the non-decreasing sequence of its associated roots.
@@
(Code Construction) Let $\ell=2$, $n_{q}\in\mathbb{N}$, and $\kappa_{1}$, $\kappa_{2},\cdots,\kappa_{n_{q}}$ be even numbers such that $|\kappa_{j}-\kappa_{j^{\prime}}|\leq 2,j,j^{\prime}\in[n_{q}]$. Then, there exists a code $\mathcal{C}$ with transition count at position j equal to $\kappa_{j},j\in[n_{q}]$ satisfying properties 1, 2, 3, and 5 in Proposition 4 such that $|\mathcal{C}|=\min\{2^{n_{q}},\sum_{j=1}^{n_{q}}\kappa_{j}\}$. Particularly, if $\kappa_{j}=2^{\delta},j\in[n_{q}]$, then there exists $\mathcal{C}$ with $|\mathcal{C}|=\min\{2^{n_{q}},n_{q}2^{\delta}\}$ satisfying properties 1-5 in Proposition 4.
The proof of Propositions 5 and 6 are provided in Appendix C and Appendix D, respectively. Using Propositions 5 and 6, we characterize the channel capacity for $\ell=2$ and $\delta>1$. Let us define $\gamma^{\prime}\triangleq\min(2^{n_{q}},n_{q}2^{\delta})$ and the set $\mathcal{T}_{\gamma^{\prime}}\subseteq\mathbb{R}^{\gamma^{\prime}}$ as the set of sequences of increasing real numbers $r_{1},r_{2},\cdots,r_{\gamma^{\prime}}$ for which there exists a partition $\{\mathcal{I}_{j},j\in[n_{q}]\}$ of $[\delta]$ such that $(r_{i_{k}},k\in\mathcal{I}_{j})$ is fully-symmetric for all $j\in[n_{q}]$, and there exists a code satisfying Properties 1-5 in Proposition 4 whose transition sets are equal to $\mathcal{I}_{j},j\in[n_{q}]$ and which has exactly one repeated codeword, i.e., only the first and last codewords are repeated. The following theorem characterizes the channel capacity.
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\in\mathbb{R}$, $\delta>1$, and let $\ell=2$. Then, the capacity is given by:
$C_{env}(P_{T},P_{ADC},h,\delta)=\max_{\mathbf{x}\in\mathbb{R}^{\gamma^{\prime}+1}}\max_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\max_{\mathbf{t}\in\mathcal{T}_{\gamma^{\prime}}}I(X;\widehat{Y}),$ (11)
where $n_{q}=\lfloor\frac{P_{ADC}}{2\alpha}\rfloor$, $\widehat{Y}=q(hX+N)$, $N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of distributions on $\{x_{1},x_{2},\cdots,x_{\gamma^{\prime}+1}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,\gamma^{\prime}-1\}$ and $q(y)=0$ if $y>t_{\gamma^{\prime}-1}$ or $y<t_{1}$.
To provide insights into the implications of the above theorem, let us consider SISO systems where the ADC thresholds are restricted to be fully-symmetric. Then, the theorem implies that the effect of using $\delta$ concatenated envelope detectors is (almost) equivalent to increasing the number of ADCs exponentially from $n_{q}$ to $2^{\delta}n_{q}$. The proof follows by similar arguments as in the proof of Theorem 1. The converse follows from Proposition 4 Item 4). Achievability follows from Proposition 6.
The region given in Theorem 2 is difficult to analyze since finding the set $\mathcal{T}_{\gamma^{\prime}}$ may be computationally complex. Inner bounds to the achievable region may be numerically derived by assuming additional symmetry restriction such as uniform quantization restrictions. This is studied in more detail in the numerical evaluations provided in Section VII.
For scenarios with $\ell>2$ and $\delta>1$, let us define $\gamma_{\ell}\triangleq\min(\ell^{n_{q}},(\ell-1)n_{q}\ell^{\delta})$ and the set $\mathcal{T}_{\gamma_{\ell}}\subseteq\mathbb{R}^{\gamma_{\ell}}$ as the set of sequences of increasing real numbers $r_{1},r_{2},\cdots,r_{\gamma_{\ell}}$ satisfying the properties in Proposition 3, for which there exists a code $\mathcal{C}$ satisfying Properties 1-5 in Proposition 4 such that i) the sets $\mathcal{P}_{j,k},j\in[n_{q}],k\in[\ell-1]$ in Proposition 3 are the indices of the codewords of $\mathcal{C}$ which have transition to or from value $k$ in their $j$th element, and ii) $\mathcal{C}$ has exactly one repeated codeword, i.e., only the first and last codewords are repeated. The following theorem characterizes the channel capacity. The proof follows from Propositions 4 and 6 similar to the arguments given in the proof of Theorem 1.
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\in\mathbb{R}$, and $\delta\in\mathbb{N}$. Then,
$\!\!\!C_{env}(P_{T},P_{ADC},h,\delta)=\max_{\begin{subarray}{c}(n_{q},\ell):\\ \alpha\ell n_{q}\leq P_{ADC}\end{subarray}}\max_{\mathbf{x}\in\mathbb{R}^{\gamma_{\ell}+1}}\max_{P_{X}\in\mathcal{P}_{\mathbf{x}}}\max_{\mathbf{t}\in\mathcal{T}_{\gamma_{\ell}}}I(X;\widehat{Y}),$ (12)
where $\widehat{Y}=q(hX+N),N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of distributions on the alphabet $\{x_{1},x_{2},\cdots,x_{\gamma_{\ell}+1}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in[\gamma_{\ell}-1]$ and $q(y)=0$ if $y>t_{\gamma_{\ell}-1}$ or $y<t_{1}$.
@@ V SISO Systems with Polynomial Operators
In this section, we evaluate the channel capacity when polynomial operators are used instead of envelope detectors.
V-A Analog Polynomial Operators
We restrict the set of implementable nonlinear analog functions to the set of polynomials of degree at most $\delta$, that is:
$\mathcal{F}^{\delta}_{poly}\triangleq\{f(y)=\sum_{i=0}^{\delta}{b_{i}}y^{i},y\in\mathbb{R}|b_{i}\in\mathbb{R},i\in[\delta]\},\quad\delta\in\mathbb{N},$
For $P_{T},P_{ADC}>0$ and $h\neq 0$, the resulting channel capacity is represented by $C_{poly}(P_{T},P_{ADC},{h},\delta)$. Similar to the case with envelope detectors, here the limitation on the degree of the polynomial functions is determined by limitations in the corresponding circuits design and is studied in Section VIII. We define the quatnizer and associate code in a similar manner as in the previous section. This is clarified through the following example.
@@ V-A Analog Polynomial Operators
Let $n_{q}=\delta=2$ and $\ell=3$ and consider a quantizer characterized by polynomials $f_{1}(y)=y^{2}+2y$ and $f_{2}(y)=y^{2}+3y,y\in\mathbb{R}$, and thresholds
$t_{1,1}=0,\quad t_{1,2}=3,\quad t_{2,1}=10,\quad t_{2,2}=18,$
@@
The following provides the SISO capacity with few-bit ADCs and polynomial operators. The proof is provided in [37].
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\neq 0$. For even-valued $\delta$, we have:
$C_{poly}(P_{T},P_{ADC},h,\delta)=\max_{\begin{subarray}{c}(n_{q},\ell):\\ \alpha\ell n_{q}\leq P_{ADC}\end{subarray}}\max_{\mathbf{x}\in\mathbb{R}^{\gamma+1}}\max_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\max_{\mathbf{t}\in\mathbb{R}^{\gamma}}I(X;\widehat{Y}),$ (13)
@@
$\max_{(n_{q},\ell):\alpha\ell n_{q}\leq P_{ADC}}\max_{\mathbf{x}\in\mathbb{R}^{\gamma+1}}\max_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\max_{\mathbf{t}\in\mathbb{R}^{\gamma}}I(X;\widehat{Y})\leq C_{poly}(P_{T},P_{ADC},h,\delta)$ (14)
$\qquad\qquad\leq\max_{(n_{q},\ell):\alpha\ell n_{q}\leq P_{ADC}}\max_{\mathbf{x}\in\mathbb{R}^{\gamma^{\prime}+1}}\max_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\max_{\mathbf{t}\in\mathbb{R}^{\gamma^{\prime}}}I(X;\widehat{Y^{\prime}}),$
where $n_{q}=\lfloor\frac{P_{ADC}}{2\alpha}\rfloor$, $\gamma\triangleq\min(\ell^{n_{q}},(\ell-1)\delta n_{q})$, $\widehat{Y}=q(hX+N)$, $N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of distributions on $\{x_{1},x_{2},\cdots,x_{\gamma+1}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,\gamma-1]$ and $q(y)=0$ if $y>t_{\gamma-1}$ or $y<t_{1}$. Furthermore we have defined $\gamma^{\prime}\triangleq\min(\ell^{n_{q}}-1,(\ell-1)\delta n_{q})$, $\widehat{Y^{\prime}}=q^{\prime}(hX+N)$, and $q^{\prime}(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,\gamma^{\prime}-1]$ and $q(y)=0$ if $y<t_{1}$ and $q(y)=\gamma^{\prime}$ if $y>t_{\gamma^{\prime}-1}$.
@@
[…]
2) The achievable region of Theorem 4 strictly subsumes Theorem 3 for equal $\gamma$ since envelope detectors generate absolute value functions which force a symmetric structure on the Voronoi regions of $q(\cdot)$. This manifests in the fully-symmetric condition $\mathbf{t}\in\mathcal{T}_{\gamma_{\ell}}$ in Theorem 3; whereas for polynomial functions, no such symmetry is required.
3) One potential approach to improve the capacity in Theorem 3 is to augment the envelope detectors by linearly combining their output with the original signal. That is, to generate operators of the form $f(y)=|y-a|+by,a,b\in\mathbb{R}$ instead of $f(y)=|y-b|,b\in\mathbb{R}$. This removes the fully-symmetric condition $\mathbf{t}\in\mathcal{T}_{\gamma_{\ell}}$. However, such linear combinations are challenging to implement using analog circuits due to timing issues in synchronizing the output of the envelope detector with the original signal.